OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. y = 4x - 2 y = x + 13

OpenStudy (anonymous):

set two equation equal to each other and solve

OpenStudy (anonymous):

(5,18)

OpenStudy (anonymous):

x + 2y = 12 -x = -y - 6

OpenStudy (anonymous):

(8,2)

OpenStudy (anonymous):

x + 2y = 12 y = x + 3